# A Joint Learning Framework for Bridging Defect Prediction and Interpretation

### v1

@@ III-C Joint learning framework
The architecture of the joint learning framework is presented in Fig. 4, each feedback loop corresponds to a specific loss function. Training stops when the model reaches the maximum number of iterations or meets the early stopping condition. Given training data set $D=\{\left(x_{i},y_{i}\right)_{i=1}^{N}\}$, $x_{i}\in R^{n}$, $y_{i}\in\{0,1\}$, the predictor $f$ and interpreter $g$, the loss function of joint learning framework is formalized as follows:
@@
where $\alpha$ and $\beta$ are hyperparameters.
The original loss function of KD consists of two essential parts: the distillation loss $L_{soft}$ (which is equivalent to $L_{int}(f,g,D)$ in Eq. (2)) and the student loss $L_{hard}$ (which is equivalent to $L_{pred}(g,D)$ in Eq. (2))). $L_{soft}$ quantifies the difference between the teacher model and the student model, which serves as a feedback loop for transferring knowledge from the teacher model to the student model. Minimizing $L_{soft}$ will enforce the student model to adopt similar decision logic to the teacher model. $L_{hard}$ evaluates the difference between the student model’s predictions and the ground truth. Minimizing $L_{hard}$ ensures that the student model retains the appropriate decision logic of the teacher model. In this paper, we introduce two key modifications to the existing loss function: (1) added the output fidelity loss $L_{pred}(f,D)$, and (2) revised the $L_{soft}$.
$L_{pred}(f,D)$ shares a similar definition with $L_{pred}(g,D)$ indicating the consistency of the predictor’s output to the ground truth.
@@
We introduce output fidelity loss guided by two critical considerations: (1) Optimizing the predictor and interpreter simultaneously. Unlike the original KD framework, which keeps the parameters of teacher model (predictor) fixed during the distillation process, minimizing the output fidelity loss will enforce the predictor to capture the discriminative features of the input metrics. (2) Improving the fidelity of the interpreter. Within the KD framework, predictor and interpreter are mutually dependent. Improving the accuracy of the predictor is equivalent to indirectly enforcing the interpreter to provide a more accurate explanation of the predictor’s decision logic. Therefore, introducing the output fidelity loss can boost the performance of the predictor and provide a more robust and insightful knowledge based on the KD principle.
Moreover, we revised the loss function $L_{soft}$ (denoted as $L_{int}(f,g,D)$ in Eq. (2)). According to Eq. (5), the $L_{int}(f,g,D)$ involves two parts, $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$. $L_{of}(f,g,D)$ is equivalent to the distillation loss in the original KD framework. The feature fidelity loss $L_{fef}(f,g,D)$ evaluates the similarity between the feature maps generated by predictor and interpreter. Minimizing $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$ can enforce the interpreter to adopt similar decision logic and feature engineering power of predictor.
@@
where $\lambda$ and $\gamma$ are hyperparameters. $L_{of}(f,g,D)$ shares a similar definition with $L_{soft}(f,g,D)$, which ensures the output of $g$ should be similar to that of $f$.
@@
To further clarify the proposed joint learning framework, we take the PC2 dataset as an example. The data in PC2 involves software metrics and labels. Software metrics serve as the input for both the predictor $f$ and interpreter $g$. Labels indicate whether a code fragment is buggy or clean. The predictor and interpreter were optimized jointly using the loss function defined in Eq. (2). Through the joint training process, the predictor was continuously refined to improve its classification accuracy. This process enables the predictor to effectively distinguish between buggy and clean code fragments. Simultaneously, the interpreter was trained to generate a decision tree that aligns with the predictor’s decision-making logic. Through this process, the predictor was also influenced by the interpreter, as minimizing $L_{int}(f,g,D)$ required the predictor to refine its feature extraction and decision logic to better align with the explanations provided by the interpreter. In Fig. 5, the decision tree generated by the interpreter reflects the optimized decision logic and highlights key decision paths for defect prediction.
Figure 5: The interpretations of the proposed method on PC2 dataset.
@@ III-D Local interpretation
It is obvious that SDT obtains the global interpretability as decision tree does [41], and interpretation for a dataset can be abstracted as a set of “IF-THEN” rules. In this section, we focus on the interpretability for an instance. Given a metric set $A=\{a_{1},a_{2},...a_{d}\}$, the sensitivity of SDT for an instance $x$ on the $a_{i}$ is defined as follows.
@@
where $N$ is the number of training data, and $a_{j}^{i}$ is the $i$-the metric of $x_{j}$. According to Eq. (8), we interpret the correlation between the instance and predictive result as the sensitivity of feature perturbation. Given an instance $x_{i}$, if a metric has a greater sensitivity, that means such a metric has a stronger influence on the prediction result.
Since the $\Delta a_{i}$ is a constant and the parameters of interpreter are fixed after training, the sensitivity $S$ can provide a stable interpretation for any instance. The algorithm for local interpretation is presented below:
@@
2:   Calculate $\text{Calculate }\bar{a}_{i}$ Eq. (10)
3:   Calculate $\text{Calculate }\Delta a_{i}$ Eq. (9)
4:   Calculate $\text{Calculate }\Delta g_{i}(x)$
@@
7:  $SI\leftarrow$ Sorts $S$ in descending order
8:  return $SI$
The loop structure (lines 2 to 5 in the algorithm) calculates the sensitivity of each metric and generates the sensitivity set $S$. Line 7 rearranges the sensitivity set $S$ in descending order and puts the index of $S$ to $SI$.
@@ IV Experiment
To objectively evaluate the performance of our method, we conducted an experiment to quantify the performance of the proposed method and the baselines on 20 datasets, aiming to answer the following research questions:
RQ1: What reliability-related improvements does the proposed method make in terms of consistency and faithfulness?
As mentioned in Section II, many local methods often yield inconsistent interpretations by adopting different data sampling techniques as well as interpreting the same code fragment multiple times. These two factors greatly affect the reliability of explainable SDP. To evaluate the consistency of interpretation, we adopt the Coincidence Degree (CD) in RQ1 [42]. Given an interpreter $g$, the CD of an instance $x_{i}$ is defined as follows :
@@
Furthermore, as discussed in Section II, another issue related to the reliability of interpretation is that predictor and interpreter may generate different predictive results for the same instance. To evaluate the consistency between predictor and interpreter, we introduce Fidelity of Interpretation (FI) [42, 43, 44] here. Given dataset $D$, predictor $f$ and interpreter $g$, the fidelity is defined as:
@@
In addition to the FI, we argue that the prediction produced by an interpreter should also align with the ground truth. This alignment indicates how much correct decision knowledge is obtained from the predictive model. In this paper, we use the Accuracy of Interpreter (AI) as another indicator to measure this alignment. Given the dataset $D$ and interpreter $g$, we define the accuracy as:
@@
RQ2: What predictive accuracy-related improvements does the proposed method make?
There are two goals in validating the predictor trained by our framework (denoted as Joint_CNN): (1) The joint learning framework has had a positive impact on the multiple evaluation metrics of the predictor, and (2) its predictive performance is comparable to other notable predictors. For the first goal, we compare the F-measure and AUC of Joint_CNN with the DP-CNN trained not by our framework (denoted as Base_CNN). In addition, we incorporate three additional metrics from [58]: Percent of Perfect Cleans (PPC), Percent of Non-Perfect Cleans (PNPC), and False Omission Rate (FOR). These indices offer valuable insights into the framework’s ability to balance classification accuracy with practical project management benefits, such as budget savings, remaining time estimates, and failure rate predictions. For the second goal, we compare the F-measure and AUC of Joint_CNN with SVM, Random Forest (RF), and Deep Belief Network (DBN) based predictors. Additionally, although SVM, RF, and DBN are commonly used for classification tasks, they are not interpretable prediction models. To better demonstrate the effectiveness and advantages of our model, we also included AutoSpearman [13], a widely recognized interpretable prediction model, in the comparison.
The Percent of Pure Cleanings (PPC) measure is defined in [58] as the percentage of correctly predicted negative cases, which indicates the amount of saved budget in the target project. PPC is calculated as:
$PPC=\frac{\left|TN\right|}{\left|nt\right|}$ (14)
where TN represents the number of true negatives and $\left|nt\right|$ denotes the total number of test instances.
The Percent of Non-Perfect Cleans (PNPC) measure is defined in [58] as the percentage representing the sum of true positives, false positives, and false negatives over the total number of instances tested. This measure is defined as follows:
$PNPC=\frac{\left|TP\right|+\left|FP\right|+\left|FN\right|}{\left|nt\right|}$ (15)
where $\left|TP\right|$ represents the number of true positives, $\left|FP\right|$ denotes the number of false positives, and $\left|FN\right|$ indicates the number of false negatives.
The False Omission Rate (FOR) measure is defined in [58] as the ratio of false negatives over the number of predicted cleans. This is used to estimate the percentage of potential failures that may occur in the target software. FOR is defined as:
@@
where $\left|FN\right|$ represents the number of false negatives, and $\left|TN\right|+\left|FN\right|$ is the total number of predicted cleans.
The F-measure, also referred to as the F1-score, assesses the balance between precision and recall. It is the harmonic mean of these two values. The F-measure is defined as:
$\color[rgb]{0,0,1}F\text{-measure}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (17)
The Area Under the Curve (AUC) is used to evaluate the performance of binary classification models by measuring the area under its Receiver Operating Characteristic (ROC) curve, which plots the trade-off between true positive rate and false positive rate across different thresholds. The AUC measures the effectiveness of a classification model in distinguishing between defective and non-defective instances. A higher AUC indicates better discrimination, where a value closer to 1 reflects excellent predictive performance.
RQ3: What is the global interpretability of the proposed method?
The objectives of RQ3 are threefold: (1) to demonstrate that our approach obtains global interpretability, (2) to verify its superior computational efficiency, and (3) to demonstrate its reliability. Although local interpretation methods explain the decision logic for a single instance, they fail to reveal the complete structure of predictors [55]. Global interpretable methods can fill in the gaps left by local interpretable methods. For the first objective, we aim to demonstrate our approach is a decision tree-like interpretable method. The substantial computational expense is a primary disadvantage of current global interpretation methods [56]. For the second objective, we want to quantify the computational efficiency of our method and SP-Lime [55] by comparing the training time of them. The rationale for using SP-Lime as the baseline is as follows. First, SP-Lime is a representative example of perturbation-based global interpretable methods and most global explainable methods are based on the perturbation algorithm [55]. Second, SP-Lime is the global version of Lime (one of baselines of RQ1). Selecting it as baseline is helpful to maintain the consistency of the experiment.
RQ4: What is the effect of different features on the performance and interpretability of our approach?
To assess the impact of individual features on the model’s performance and interpretability, we carried out an ablation experiment. Specifically, we took Joint_CNN as the predictor and eliminated metrics individually. The impacts on the predictor was evaluated by observing changes in the F-measure, AUC, and MCC, and compared these results with the model’s interpretability. MCC is an acronym for Matthews Correlation Coefficient that is used to measure the quality of binary classifications, particularly in imbalanced datasets. The MCC is defined as follows:
$\color[rgb]{0,0,1}MCC=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (18)
where $TP$, $TN$, $FP$ and $FN$ represent the number of true positives, true negatives, false positives, and false negatives, respectively. The MCC value ranges from -1 to 1, where -1 indicates complete disagreement between predictions and actual outcomes, 1 represents perfect agreement, and 0 implies performance equivalent to random guessing.
@@ IV-A Dataset
The 20 datasets used in the experiment were collected from the AEEEM 111 https://zenodo.org/record/3362613#.YmuGoNpByUk, NASA  [^2] http://promise.site.uottawa.ca/SERepository/datasets-page.html and PROMISE  [^3] http://promise.site.uottawa.ca/SERepository repositories. The effectiveness of these datasets has been demonstrated by numerous studies [45, 46, 47, 48]. Further details about each dataset are provided in Table II. We observe that these datasets exhibit significant class imbalance, with defective instances accounting for only 0.41% of the PC2 dataset.
@@ IV-B Setting
In the experiment, some parameters were set as follows: the initial learning rate $lr$ = 1e-06 and the batch size $batch\_size$ is set to 16. The depth of SDT $tree\_depth$ is set to 4. The penalty strength of SDT $penalty\_strength$ is set to 1e+1 and the penalty decay rate $penalty\_decay$ is set to 0.25. The hyperparameters $\gamma$, $\lambda$, $\alpha$ and $\beta$ are set to 0.8, 0.5, 1.4 and 0.6 respectively. The temperature hyperparameter T is set to 100. The window size for the Exponential Moving Average (EMA) $ema\_win\_size$ is set to 1000. We applied the same set of parameters across all datasets to ensure a fair comparison. The dataset was split into 7:2:1 for training, testing, and validation. The experiments are implemented by TensorFlow 2.3.0 with the NVIDIA GeForce RTX 3090 GPU.
@@ IV-B1 Baseline
We utilize LIME [8] and BreakDown [28] as baselines for RQ1 and RQ3. We selected these techniques due to the following reasons: (1) These two methods are extensively employed by the SDP research community; (2) LIME, BreakDown, and our approach belong to feature-based interpretation techniques that strive to quantify the correlations between input features and predictive outputs; (3) The theoretical foundations of LIME and BreakDown differ substantially, thereby enriching the diversity of comparisons. LIME approximates the behavior of predictors in a local context, whereas BreakDown calculates the cumulative contribution of each feature to the prediction result.

### v2

@@ III-C Joint learning framework
The architecture of the joint learning framework is presented in Fig. 4, each feedback loop corresponds to a loss function. Given training data set $D=\{\left(x_{i},y_{i}\right)_{i=1}^{N}\}$, $x_{i}\in R^{n}$, $y_{i}\in\{0,1\}$, the predictor $f$ and interpreter $g$, the loss function for joint learning framework is defined as follows:
@@
where the hyperparameters $\alpha$ and $\beta$ denote the relative importances of the subloss to the the overall loss.
The original loss function of KD consists of two parts: the distillation loss $L_{soft}$ (which is equivalent to $L_{int}(f,g,D)$ in Eq. (2)) and the student loss $L_{hard}$ (which is equivalent to $L_{pred}(g,D)$ in Eq. (2)). $L_{soft}$ quantifies the difference between the teacher model and the student model, which serves as a feedback loop for transferring knowledge from the teacher model to the student model. $L_{hard}$ evaluates the difference between the student model’s predictions and the ground truth.
In this paper, we introduce two key modifications to the original KD loss function: (1) added the output fidelity loss $L_{pred}(f,D)$, and (2) revised the $L_{soft}$. $L_{pred}(f,D)$ shares a similar definition with $L_{pred}(g,D)$ indicating the consistency of the predictor’s output to the ground truth.
@@
Introducing output fidelity loss here can enhance both prediction accuracy and interpreter fidelity simultaneously. Within the KD framework, predictor and interpreter are mutually dependent. Minimizing the output fidelity loss will encourage the predictor to achieve heigher accuracy. Moreover, improving the accuracy of the predictor is equivalent to indirectly enforcing the interpreter to provide a more accurate explanation of the predictor’s decision logic. Therefore, incorporating the output fidelity loss can boost the performance of the predictor but also yield a more robust and insightful knowledge based on the KD principle.
The revised the loss function $L_{soft}$ (denoted as $L_{int}(f,g,D)$ in Eq. (2)) is presented in Eq. (5). It involves two parts, $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$. $L_{of}(f,g,D)$ is equivalent to the distillation loss in the original KD framework. The feature fidelity loss $L_{fef}(f,g,D)$ evaluates the similarity between the feature maps generated by predictor and interpreter. Minimizing $L_{of}(f,g,D)$ and $L_{fef}(f,g,D)$ can enforce the interpreter to achieve similar feature engineering power of predictor.
@@
where $\lambda$ and $\gamma$ are hyperparameters. $L_{of}(f,g,D)$ shares a similar definition with $L_{soft}(f,g,D)$.
@@
To further demonstrate the proposed joint learning framework, we use the PC2 dataset as an example. Each data point in PC2 consists of twenty software metrics and one label. According to Eqs. (3, 4), DP-CNN and SDT iteratively optimize their parameters to minimize discrepancies between predictions and actual labels. Crucially, improving prediction accuracy of DP-CNN inherently enforces the SDT to provide a more precise explanation of the predictor’s decision-making process, which indirectly improve interpreter’s reliability. Meanwhile, according to Eq. (6), the features generated by DP-CNN and SDT should be similar, ensuring that both models have comparable decision logic, thereby enhancing the interpretability of SDT. Additionally, Eq. (7) ensures consistency between SDT’s selected features and the input metrics, which strengthens the reliability of the interpreter. As presented in Fig. 5, the decision logic of DP-CNN is interpretered as a set of “IF-THEN” paths in the SDT, which is similar to the decision tree doess [41].
Figure 5: The interpretations of the proposed method on PC1 dataset.
@@ III-D Local interpretation
In this section, we focus on the interpretability for an instance. Given a metric set $A=\{a_{1},a_{2},...a_{d}\}$, the sensitivity of SDT for an instance $x$ on the $a_{i}$ is defined as follows.
@@
where $N$ is the number of training data, and $a_{j}^{i}$ is the $i$-th metric of $x_{j}$. According to Eq. (8), we interpret the correlation between the instance and predictive result as the sensitivity to feature perturbation. Given an instance $x_{i}$, if a metric has a greater sensitivity, this means the metric has a stronger influence on the prediction result.
Since the $\Delta a_{i}$ is a constant and the parameters of interpreter are fixed after training, the sensitivity $S$ can provide a stable and consistent interpretation for any instance. The algorithm for performing local interpretation is presented below:
@@
2:  $\text{Calculate }\bar{a}_{i}$ Eq. (10)
3:  $\text{Calculate }\Delta a_{i}$ Eq. (9)
4:  $\text{Calculate }\Delta g_{i}(x)$
@@
[…]
8:return SISI
The loop structure (lines 2 to 5 in the algorithm) computes the sensitivity of each metric and generates the sensitivity set $S$. Line 7 rearranges the sensitivity set $S$ in descending order and puts the index of $S$ to $SI$.
@@ IV Experiment
To objectively evaluate the performance of our method, we conducted an experiment to quantify the performance of the proposed method and the baselines on twenty datasets, aiming to answer the following research questions:
RQ1: What improvements does the proposed method offer for the consistency of interpretations?
As mentioned in Section II, many local methods tend to yield inconsistent interpretations, which greatly affects the reliability of explainable SDP. To evaluate the consistency of interpretation, we adopt the Coincidence Degree (CD) in RQ1 [42]. Given an interpreter $g$, the CD of an instance $x_{i}$ is defined as follows :
@@
Furthermore, another issue related to the consistency of interpretation is that the predictor and interpreter may generate different predictive results for the same instance. To evaluate the consistency between predictor and interpreter, we introduce Fidelity of Interpretation (FI) [42, 43, 44] here. Given dataset $D$, predictor $f$ and interpreter $g$, the fidelity is defined as:
@@
In addition to FI, we also argue that the prediction produced by an interpreter should also align with the ground truth. The alignment indicates how much correct decision knowledge is captured from the predictor. In this paper, we use the Accuracy of Interpreter (AI) to measure this alignment. Given the dataset $D$ and interpreter $g$, we define the accuracy as:
@@
RQ2: What performance improvements does the joint-learning framework achieve?
There are two goals in validating the predictor trained by the joint learning framework (denoted as Joint_CNN): (1) to assess the positive impact of the joint learning framework on predictor’s performance, and (2) to determine whether its predictive performance outperforms other notable predictors. For the first goal, we evaluate model performance using five key indicators: F-measure, AUC, Percent of Perfect Cleans (PPC), Percent of Non-Perfect Cleans (PNPC), and False Omission Rate (FOR). Although F-measure and AUC provide an overall assessment of predictive performance, they can not fully capture the effectiveness in practical SDP scenarios. To address this gap, we introduce PPC, PNPC, and FOR. For the second goal, we compare the F-measure and AUC of Joint_CNN with existing explainable and non-explainable SDP models. PPC, PNPC and FOR are defined as follows:
$PPC=\frac{\left|TN\right|}{\left|D\right|}$ (14)
$PNPC=\frac{\left|TP\right|+\left|FP\right|+\left|FN\right|}{\left|D\right|}$ (15)
@@
where $\left|TP\right|,\left|FP\right|,\left|FN\right|,\left|FN\right|$ represent the number of true positives, false positives, true negative and false negatives.
The F-measure assesses the balance between precision and recall.
$F\text{-measure}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (17)
The AUC is used to evaluate the performance of binary classification models by measuring the area under its Receiver Operating Characteristic (ROC) curve.
RQ3: what’s the performance of our method in terms of global interpretability?
The RQ3 involves three objectives: (1) evaluating the global interpretability of our approach, (2) assessing its computational efficiency, and (3) demonstrating its reliability. For the first objective, we aim to demonstrate that our approach provides decision tree-like interpretability. The decision tree is a typical global interpretation method that represents the overall decision logic through hierarchical “IF-Then” rules [45]. Furthermore, we adopt the same quantitative method as [46] to evaluate global interpretability by measuring the contributions of metric subsets, identified by the interpreter, to the model’s predictive performance. Specifically, we use Joint_SDT and SP-LIME to generate six metric subsets for each dataset and use them to retrain Joint_CNN independently. The predictive performance of Joint_CNN with different metric subsets is measured by AUC. Higher AUC values indicate that the interpreter provides better global interpretability. For the second objective, we assess the computational efficiency of our method and SP-Lime by comparing the training times. Since excessive training time is a key limitation of existing global interpretation methods [47]. For the third objective, we use three indicators, CD, FI and AI to quantify the reliability of our approach.
RQ4: Do the metrics selected by Joint_SDT have a significant impact on the performance of Joint_CNN?
This study evaluates the consistency between the metrics selected by Joint_SDT and those impact the performance of Joint_CNN most. To address this, we conduct ablation experiments in which each metric identified by Join_SDT is individually removed, and the resulting degradation in Joint_CNN’s performance (e.g., declines in Matthews Correlation Coefficient (MCC), F-measure, and AUC) is measured. MCC is commonly used to assess the quality of imbalanced binary classification, where the value ranges from -1 to 1. -1 indicates complete disagreement between predictions and ground truth, 1 represents perfect agreement. A strong correlation between Joint_SDT’s prioritized metrics and the magnitude of performance deterioration would further confirm that our joint-learning framework make positive contributions to both predictor and interpreter.
$MCC=\frac{TP\cdot TN-FP\cdot FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$ (18)
(a) The first interpretation
(b) The second interpretation
(c) The third interpretation
(d) The fourth interpretation
@@ IV-A Dataset
Twenty datasets used in the experiment are collected from the AEEEM 111 https://zenodo.org/record/3362613#.YmuGoNpByUk, NASA  [^2] http://promise.site.uottawa.ca/SERepository/datasets-page.html and PROMISE  [^3] http://promise.site.uottawa.ca/SERepository repositories. Numerous studies have demonstrated the efficacy of these datasets [48, 49, 50, 51]. Further details about each dataset are provided in Table II. We observe that these datasets exhibit significant class imbalance, with the highest imbalance rate reaching 0.41% from the PC2 dataset.
@@ IV-B Setting
We config our model as follows: the learning rate is initialized as 1e-06 and the batch size is set to 16. SGD optimizer is employed with a momentum of 0.9. The depth of SDT is set to 4. The penalty strength of SDT is set to 1e+1 and the penalty decay rate is set to 0.25. The hyperparameters $\alpha$ and $\beta$ in Eq. (2) are set to 1.4 and 0.6, and $\gamma$, $\lambda$ in Eq. (2) are set to 0.8 and 0.5. The temperature hyperparameter T in Eq. (6) is set to 100. The window size for the Exponential Moving Average (EMA) is set to 1000. Training stops when the model reaches the maximum number of iterations (40) or when early stopping criterias are met (no validation loss improvement for 10 consecutive epochs). Based on preliminary experimental results, the maximum number of iterations is set to 40, beyond which model performance saturates with minimal further improvement. The dataset was split into 7:2:1 for training, testing, and validation. The experiments are implemented by TensorFlow 2.3.0 with the NVIDIA GeForce RTX 3090 GPU. To ensure fairness, we applied this configuration to the experiments on all datasets.
@@ IV-B1 Baseline
For RQ1, we utilize LIME [12] and BreakDown [30] as baselines for the following reasons: (1) These methods are extensively used in the SDP research; (2) LIME, BreakDown, and our approach belong to feature-oriented interpretation techniques that aim to quantify the correlations between input features and predictive outputs; (3) The theoretical foundations of LIME and BreakDown are different, thereby enriching the diversity of comparisons. For RQ2, we use SVM, Random Forest (RF), Deep Belief Network (DBN), and AutoSpearman [52] as baselines. These models are selected based on the following reasons: (1) SVM and RF are widely adopted non-explainable SDP models, while DBN represents DL-based approach in SDP. (3) AutoSpearman, in contrast, is a well-recognized explainable SDP model. Including it as a baseline ensures a fair comparison. For RQ3, we use SP-LIME [45] as the baseline for two key reasons: (1) SP-LIME and our method share the same technical foundation. Both of them are perturbation-based global interpretable methods. Moreover, SP-LIME is the representative method in this category [45]. (2) SP-LIME is the global version of Lime (one of the baselines of RQ1). Selecting it as the baseline is helpful to maintain the consistency across the experiments.